# Predicting and Understanding Turn-Taking Behavior in Open-Ended Group Activities in Virtual Reality

### v1

@@ 2.3. Social Interactions across Individuals and Groups
The ability of tracking data to predict individual differences such as personalities and affective states (Wampfler et al., 2022; Kovačević et al., 2023; Kampman et al., 2018; Ivanov et al., 2011) suggests that individual differences can be used to also predict user behaviors. Individual differences such as gender, age, and personalities were shown to be related to how people take up physical and virtual spaces (Friebel et al., 2021; Iachini et al., 2016; Nassiri et al., 2010; Iachini et al., 2014), while personality, public speaking anxiety, and immersive tendencies of individuals can be predictive of their experience and perception of the quality of interactive systems (Chollet et al., 2018).
Personality traits, in particular, are also related to speech behaviors. For example, the personality traits of conscientiousness and extraversion are related to basic speech features such as pitch estimate (Ivanov et al., 2011). The levels of extraversion and neuroticism are also associated with different speech patterns, notably with introverts exhibiting longer silences between speech utterances (Ramsay, 1968). Focusing on the extraversion trait, introverts and extroverts were found to speak at different levels of language abstraction (Beukeboom et al., 2013) and also differ in reaction time when verbally responding to prompts and questionnaire items (Park et al., 2020; Lee et al., 2021). Broadly speaking, personality traits of extraversion, agreeableness, and neuroticism are further correlated with varying features of speaking turns related to speaking turns, speaking length, and average speaking turn duration (Aran and Gatica-Perez, 2013; Lepri et al., 2010).
Finally, social interactions can vary depending on characteristics of social groups. For example, the association between trait dominance – measured as expressed control and prosocial interpersonal power and influence – and speaking time is influenced by the composition of the social groups, namely whether the group was composed randomly or of extremes (i.e., pairing together most and least dominant individuals) (Mast, 2002). Comparing conversations between dyad and triads, the presence of an additional listener led to individuals speaking louder in situations with high noise levels, and listeners orienting their heads more optimally (i.e., rotating their heads to approximately 30 degrees from the speaker) (Hadley et al., 2021).
We rely on past findings to identify and extract features related to personality and group compositions, and used them to model turn-taking behaviors. By studying how these features are related to social behaviors such as speaking intentions, we explore how incorporations of individual- and group- level characteristics can help build immersive platforms and tools for facilitating more personalized and adaptive social interactions.
![Refer to caption](extracted/5634250/figures/engage-v2.png)
@@ 3.1. Recordings of Open-ended Group Social Interactions in VR
We studied the large-scale longitudinal VR classroom dataset collected by Han et al. (Han and Bailenson, 2024), and focused on the subset of the dataset that was collected in Fall 2022 as part of a university course on VR and its application. In it, 146 university students, out of whom 117 consented, engaged in weekly group discussions and responded to open-ended design prompts in groups of size two to four for four weeks. A research personnel not part of the teaching staff worked to randomly assign students into groups (fixed over time), and at the same time maximize the number of groups made up with all consented students. This approach ensured that students received comparable learning experience as the teaching staff was blinded to the consenting status of the group, while also allowed researchers to later work with large amounts of VR social interactions. Given our focus on turn taking, we only analyzed VR sessions with all consenting participants.
@@
The topics of discussion and activities changed weekly, with themes such as accessibility and education. Each week, students gathered in the same preassigned groups using the social VR platform ENGAGE, while being physically remote in their own spaces (e.g., dorm room, office space). Each student used a Meta Quest 2 headset and two hand-held controllers to partake in the sessions as they were allowed to move around the virtual spaces using smooth translation and teleportation. No teaching staff was present in these group sessions, so student groups recorded the discussion sections using the platform’s built-in recording feature. Figure 1 shows screenshots from the recorded open-ended group activities.
@@
The authors varied the characteristics of the virtual environments (i.e., ceiling height, amount of visible space, indoor vs. outdoor) to examine how context influences attitude, nonverbal behavior, and design behaviors (Han et al., 2024; Wang et al., 2024a). This variation in virtual context allows us to model interactions occurring in diverse virtual spaces, setting it apart from prior literature that examined speech behavior in the same physical or virtual environment (Chen et al., 2024; Lee et al., 2023). Another reason for choosing this particular dataset was that the discussion sections were based on open-ended group activities, not instructor-led, and therefore consisted of more natural group social interactions compared to the rest of the VR classroom dataset. While previous research have examined language use (i.e., linguistic style from session-level verbal transcripts) using the subset of data collected in Fall 2021 (DeVeaux et al., ; DeVeaux et al., 2023), our work is the first to use the Fall 2022 data for examining turn-taking behaviors of open-ended group discussions.
@@
As we were interested in understanding group social interactions, we further filtered out moments and sessions with only two students. For predicting turn-taking behaviors such as who the next speaker is, looking at groups larger than two also makes the problem nontrivial. Our final dataset consisted of 77 VR sessions collected from 26 groups and 100 unique students. Some sessions were dropped due to technical difficulties such as software updates and incomplete recordings. The dataset aggregated a total of 1660.60 minutes of tracking data, with each session taking on average 21.57 minutes (SD = 6.44). We provide additional details on the group sessions in Appendix A.
@@
The recording files collected several forms of user- and session- level data, namely that of the scene (e.g., user inserted objects, drawings), user motion, and audio. Since we wanted to extrapolate turn-taking behavior insights that are generalizable across social platforms and settings (e.g., different virtual environments and activities), we focused on utilizing the information related to user motion and audio, which is logged at 30 Hz. User motion data consisted of the position (i.e., x, y, z) and orientation (i.e., roll, pitch, yaw) of the user’s headset, two controllers, and the “root”. The “root” tracks the user’s global position and orientation with respect to the virtual environment, while the headset and controllers are tracked with respect to the coordinate system dictated by the “root”. Using the tracked user motion, we were able to derive both information regarding the user’s egocentric behavior, for example where their right hand is positioned with respect to their body, as well as those related to other users, such as the interpersonal distance and eye gaze. The audio information is recorded both as a global audio file as well as a floating-point value between 0 and 1 that represents each user’s speech volume.
@@ 3.2. Categories of Turn-Transition Behaviors
To investigate social behaviors, we began by identifying all speech events by labeling Inter-Pausal-Units (IPU), defined as the stretches of speech activity by a single speaker. We set the threshold for speech volume of determining speech activity to 0.1. To establish a more consistent labeling of speakers during short pauses during speaking activities, we applied smoothing (i.e., joining adjacent speech events of the same user) to the raw labels when the gap between adjacent speech events is within 0.5s. We chose the threshold of 0.5s based on prior literature that found the average pause duration during read speech, interviews and public presentations to be between 0.38 to 0.53 seconds (Liu et al., 2022; Campione and Véronis, 2002).
@@
For each session recording, we then proceeded to determine the main speaker by first labeling users who are the sole speaker at a certain time frame as the main speaker. When multiple users are speaking at the same time, we implemented the following labeling scheme for determining the main speaker. To start, we eliminated speech events that are completely overlapped by another speech event. Then, we assign the main speaker as the speaker who ends their speech event last. The start time of the new main speaker is marked as the moment the previous main speaker finished speaking.
@@
From the labeled main speakers, we extracted four types of turn-transition behaviors following closely the categorizations used by Jokinen et al. (Jokinen et al., 2013), namely (1) clean turn taking, (2) overlap turn taking, (3) backchanneling, and (4) continuing speech. These categories provide a foundation for our examination of turn-taking behaviors, which placed a greater focus on the clean turn taking and overlap turn taking categories. Figure 2 shows categorization of each of the categories based on an example raw audio input. Upon obtaining the labels, we removed data points associated with speech events shorter than 323 milliseconds (i.e., the average duration for enunciating fast words found in (Sommers and Barcroft, 2006)) to filter out noise.
![Refer to caption](extracted/5634250/figures/catergorization-v3.png)
@@
Clean Turn Taking. Clean turn taking occurs whenever the main speaker has changed and that there was no overlapping between the previous and new speaker turns. The start of a clean turn taking event was marked as the beginning of the new speaker’s speech event. Figure 2b-2 shows examples of these behaviors.
@@
Overlap Turn Taking. As shown in Figure 2b-3, overlap turn taking occurs whenever the main speaker changes and the two speech events overlap. In other words, the start of the speech event of the new speaker precedes the end of the previous speech event. We again marked the start of a turn taking event as the beginning of the new speaker’s speech event.
Backchanneling. We determined backchanneling by checking for instances when a non-main speaker began a speech event after another speaker began their speech event and stopped speaking before the other speaker ended their speech event. The start of a complete overlap event is labeled as the start of the speech event of the non-main speaker. We show examples of backchanneling in Figure 2b-4.
@@
Continuing Speech. Continuing speech is similar to the scattered contribution category proposed by Jokinen et al. (Jokinen et al., 2013). In their formulation, scattered contributions are defined as two adjacent speech instances of the same speaker being spaced less than 200ms apart, while those that are further apart are considered sequences of turns by the same speaker. Since we are interested in understanding how and when turn transitions take place (i.e., changes in speakers), we combined these two scenarios under the broader category of continuing speech. More concretely, we label continuing speech by first locating instances of adjacent speech events of the same user and then marking the start of these turn-transition behaviors as the beginning of the later speech event. Figure 2b-5 shows instances of continuing speech behaviors.
@@ 3.3. Feature Selection
We formulated features from the tracking data and user demographics based on prior literature in social dynamics. As reviewed in Sections 2.1 and 2.2, user motion and verbal behaviors are predictive of user intentions, identity and social dynamics. Past works have also shown that information relating to the current speakers (Chen et al., 2024) as well as the dyadic relationship between the current speakers and listeners (Ishii et al., 2016b, 2017) can offer insights to model speech behaviors.
With this in mind, we extracted features related to individual listeners and current speakers, the dyadic dynamic between the listener and speaker, their dynamics with the rest of the group, as well as verbal behaviors. Formally, we focused on encoding features related to two users: a user of interest which we refer to as the main user, and the previous speaker, which we refer to as the reference user. To extract motion-related features, we examined the tracking data across the 1-second window at moments prior to the start of turn-transition behaviors, with the specific timestamps used differing across prediction tasks. The 1-second window size is comparable to what was used in past literature on predicting speaking intentions (Chen et al., 2024) and is determined based on piloting other durations. We present a summary of the extracted features in Table 1.
@@
Dyadic Relationship (Chen et al., 2024; Jokinen et al., 2013; Ishii et al., 2016b; Lee et al., 2023) | Direct gaze from the main user to the reference user | 3 summary stats. 2 meas. = 6 |
Direct gaze from the reference user to the main user | 3 summary stats. 2 meas. = 6 | |
@@
Direct gaze from a user to all remaining users, where the user is either the main user or reference user | 3 summary stats. 2 meas. 2 users = 12 |
Direct gaze from all remaining users to a user, where the user is either the main user or reference user | 3 summary stats. 2 meas. 2 users = 12 | |
@@ 3.3.1. Speech-Related Features
The first group of features is related to speech behaviors, as they have been shown to be predictive of who the next speaker is (Parker, 1988). Following the methods of Parker (Parker, 1988), we first encoded speech sequences leading up to the start of the examined window. To do this, we begin at the timestamp at the start of the examined window, and trace backwards each time the main speaker changes, at which point we increment the turn index by 1. If the main speaker for a given turn is the main user, we assign the feature value at the corresponding turn index “u”. If we encounter a new speaker at a preceding turn, we assign the feature value a new symbol corresponding to that user (e.g., “a”, “b”, “c”). For turns with a speaker who we had already created a symbol for, we assigned their mapped symbol to the feature value. In our setup, we encoded the 10 preceding speaker indices, assigning “NA”s to features when there were no more changes in main speakers preceding the last encoded turn. In practice, we further processed the categorical speech sequence features at each turn index using one-hot encoding.
We also introduced three additional features to describe the main user’s verbal behaviors, namely (1) the number of speaking turns before the main speaker’s last speech event, the (2) the duration of time before the end of the main user’s last speech event, and (3) a binary feature encoding whether the main user has spoken before during the current session. It was unnecessary to derive these features for the reference user as they were by definition the speaker of the previous turn.
@@ 3.3.2. Individual- and Group- Level Features
Drawing from works that indicate correlations between speech behaviors and personalities and group compositions (Ramsay, 1968; Ivanov et al., 2011; Aran and Gatica-Perez, 2013; Mast, 2002; Hadley et al., 2021), we also formulated features based on individual and group characteristics. Specifically, we encoded the Big-5 personality traits (Gosling et al., 2003) of the main user and reference user, as well as the average across the entire group. In our examined dataset (Han and Bailenson, 2024), the authors measured Big-5 personality using the ten item personality measure (TIPI) introduced by Gosling et al. (Gosling et al., 2003), with each personality trait ranging between 1–7 and calculated as the mean of two of the ten items.

### v2

@@ 2.3. Social Interactions across Individuals and Groups
Tracking data’s ability to predict individual differences such as personalities and affective states (Wampfler et al., 2022; Kovačević et al., 2023; Kampman et al., 2018; Ivanov et al., 2011) suggests that individual differences can be used to predict user behaviors. Individual differences such as gender, age, and personalities are related to how people take up physical and virtual spaces (Friebel et al., 2021; Iachini et al., 2016; Nassiri et al., 2010; Iachini et al., 2014), while personality, public speaking anxiety, and immersive tendencies of individuals are predictive of user experience and perceived quality of interactive systems(Chollet et al., 2018).
Personality traits, in particular, are related to speech behaviors. For example, conscientiousness and extraversion are related to basic speech features such as pitch estimate (Ivanov et al., 2011). Extraversion and neuroticism are also associated with different speech patterns, notably with introverts exhibiting longer silences between utterances (Ramsay, 1968). Introverts and extroverts were also found to speak at different levels of language abstraction (Beukeboom et al., 2013) and differ in reaction time when verbally responding to prompts and questionnaires (Park et al., 2020; Lee et al., 2021). Broadly speaking, extraversion, agreeableness, and neuroticism also correlate with features related to speaking turns, speaking length, and average speaking turn duration (Aran and Gatica-Perez, 2013; Lepri et al., 2010).
Broadly speaking, social interactions can vary depending on the characteristics of social groups. For example, the association between trait dominance—measured as expressed control and prosocial interpersonal power and influence—and speaking time is influenced by the composition of social groups, namely whether the group was composed randomly or of extremes (i.e., pairing together the most and least dominant individuals) (Mast, 2002). Comparing conversations between dyad and triads, the presence of an additional listener led to individuals speaking louder in situations with high noise levels, and listeners orienting their heads more optimally (i.e., rotating their heads to approximately 30 degrees from the speaker) (Hadley et al., 2021).
We rely on past findings to identify features related to personality and group compositions for modeling turn taking. By studying how these features are related to social behaviors such as speaking intentions, we explore how individual and group characteristics can help build immersive tools for facilitating social interactions.
@@ 3.1. Recordings of Open-ended Group Social Interactions in VR
We studied the longitudinal VR classroom dataset collected by Han et al. (Han and Bailenson, 2024), and focused on the subset collected in Fall 2022 as part of a university course on VR. In it, 146 university students, out of whom 117 consented, engaged in weekly discussions and responded to open-ended design prompts in groups of two to four for four weeks. Given our interest in turn taking, we only analyzed VR sessions with all consenting participants111A research personnel not part of the teaching staff randomly assigned students into groups and maximized the number of groups with all consenting students. This ensured that students received comparable learning experience as the teaching staff was blinded to the consenting status, while also allowed researchers to examine fully-consenting group interactions.. The topics of discussion and activities changed weekly, with themes such as accessibility and education. We detail session activities in Appendix A. Each week, students gathered in the same groups and used the social VR platform ENGAGE while being physically remote in their own private spaces. Each student used a Meta Quest 2 headset and two hand-held controllers to partake in the sessions and was allowed to move around virtually using smooth translation and teleportation. No teaching staff was present in these sessions, so student groups recorded the discussion sections using the platform’s recording feature. Figure 1 shows screenshots from the recorded group activities.
@@
The authors varied the virtual environments (i.e., ceiling height, amount of visible space) to examine how context influences attitude, nonverbal behavior, and design behavior (Han et al., 2024; Wang et al., 2024a). This variation in context allows us to model interactions occurring in diverse virtual spaces, setting it apart from literature that examined speech behavior in the same physical or virtual environment (Chen et al., 2024; Lee et al., 2023). This particular dataset was also well-suited for our purpose given its open-ended activities, which fostered more natural interactions compared to the instructor-led sections in the rest of the dataset. While previous research have examined language use using the data collected in Fall 2021 (DeVeaux et al., ; DeVeaux et al., 2023), we are the first to use the Fall 2022 data to examine turn-taking behaviors of open-ended group activities.
@@
Given our interest in group interactions, we filtered out moments and sessions with two students. For predicting turn-taking behaviors, looking at groups larger than two also makes the problem nontrivial. Our final dataset consisted of 77 VR sessions collected from 26 groups and 100 unique students. Out of the 77 sessions, 35 were three-person discussions, and 42 were of four people. During data collection, some participants missed discussions, leading to a high number of three-person discussions. Some sessions were dropped due to technical difficulties such as software updates and incomplete recordings. The dataset aggregated a total of 1660.60 minutes of tracking data, with each session taking on average 21.57 minutes (SD=6.44).
@@
The recording files collected several forms of user and session data, namely that of the scene, user motion, and audio. Since we wanted to extrapolate turn-taking behavior insights that are generalizable across social platforms and social settings, we focused on encoding information related to user motion and audio, which is logged at 30 Hz. Motion data consisted of the position (i.e., x, y, z) and orientation (i.e., roll, pitch, yaw) of the user’s headset, two controllers, and the “root”. The “root” tracks the user’s global position and orientation within the virtual environment, while the headset and controllers are tracked within the coordinate system dictated by the “root”. From the tracked motion, we can derive information regarding a user’s egocentric behavior as well as those related to other users. The audio information is recorded both through audio files and a floating-point value between 0 and 1 representing volume.
@@ 3.2.1. Pre-processing Speech Events
We began by identifying all speech events by labeling Inter-Pausal-Units (IPU) (Koiso et al., 1998), defined as the stretches of speech activity by a single speaker. Using the audio tracking data, we determined users as actively speaking when their speech volume is greater than 0.1. To reduce noise from short pauses between speaking activities, we joined adjacent speech events of the same user when the gap between them is within 0.5s. The threshold of 0.5s follows prior literature that found the average pause duration during read speech, interviews and public presentations to be between 0.38 to 0.53 seconds (Liu et al., 2022; Campione and Véronis, 2002).
@@
For each session recording, we proceeded to determine the main speaker by first labeling users who are the sole speaker during speech events as the main speaker. When there are multiple speakers, we implemented the following labeling scheme. To start, we eliminated speech events that are completely overlapped by another speech event. Then, we assign the main speaker as the speaker who ends their speech event last. The start time of the new main speaker is marked as the moment the previous main speaker finished speaking. We used Python for extracting IPUs and assigning main speakers.
3.2.2. Defining and Labeling Turn-transition Behaviors
@@
From the labeled main speakers, we extracted four types of turn-transition behaviors following closely the categorizations used by Jokinen et al. (Jokinen et al., 2013): clean turn taking, overlap turn taking, backchanneling, and continuing speech. These categories provide a formal framework for us to examine turn-taking behaviors. Upon obtaining the labels, we removed data points associated with speech events shorter than 323 milliseconds (i.e., the average duration for enunciating fast words found in (Sommers and Barcroft, 2006)) to filter out noise. We used Python for processing the audio input. Figure 2 shows categorization of each of the categories based on an example audio input. Table 1 shows the categories’ summary statistics.222A repository with code used for labeling the four categories is available at github.com/pwang1230/turntaking-wang-2025.
[b]
Week 1 | Week 2 | Week 3 | Week 4 | Total | |
Clean Turn Taking | 1675 (4.42) | 2386 (4.71) | 1712 (5.12) | 2149 (4.87) | 7922 (4.77) |
Overlap Turn Taking | 622 (1.64) | 939 (1.86) | 771 (2.30) | 904 (2.05) | 3236 (1.95) |
Backchanneling | 571 (1.51) | 947 (1.87) | 754 (2.25) | 771 (1.75) | 3043 (1.83) |
Continuing Speech | 1866 (1.12) | 2446 (4.83) | 1600 (4.78) | 2292 (5.20) | 8204 (4.94) |
@@
Clean Turn Taking. Clean turn taking occurs whenever the main speaker has changed and that there was no overlapping between the previous and new speaker turns. The start of a clean turn taking event was marked as the beginning of the new speaker’s speech event. Figure 2b-1 shows examples of these behaviors.
@@
Overlap Turn Taking. As shown in Figure 2b-2, overlap turn taking occurs whenever the main speaker changes and the two speech events overlap. In other words, the start of the speech event of the new speaker precedes the end of the previous speech event. We again marked the start of a turn taking event as the beginning of the new speaker’s speech event.
Backchanneling. We determined backchanneling by checking for instances when a non-main speaker began and ended their speech event without ever being the main speaker. The start of a backchanneling event is labeled as the start of the speech event of the non-main speaker. We show examples of backchanneling in Figure 2b-3. One noteworthy scenario to this labeling scheme is when the non-main speaker’s speech event covers a change in the main speaker. As shown in Figure 2b-3, specifically the final backchanneling event by the bottom user, the main speaker changed during the speech event. Since the bottom user was never labeled the main speaker during their speech event, we labeled this instance as backchanneling.
@@
Continuing Speech. Continuing speech is similar to the scattered contribution category proposed by Jokinen et al. (Jokinen et al., 2013). In their formulation, scattered contributions are defined as two adjacent speech instances of the same speaker being spaced less than 200ms apart, while those that are further apart are considered sequences of turns by the same speaker. Since we are interested in understanding how and when turn transitions take place (i.e., changes in speakers), we combined these two scenarios under the broader category of continuing speech. More concretely, we label continuing speech by first locating instances of adjacent speech events of the same user and then marking the start of the later speech event as the beginning of the turn-transition behavior. Figure 2b-4 shows instances of continuing speech behaviors.
@@ 3.3. Feature Selection
We formulated features based on prior literature. As reviewed in Section 2, user motion, individual and group differences, and verbal behaviors are predictive of user intentions and social dynamics. Past works have also shown that information relating to the current speaker (Chen et al., 2024) and their relationship with listeners (Ishii et al., 2016b, 2017) offer insights to speech behaviors. Importantly, although our dataset varied the virtual spatial context, we chose to not encode these differences because they were manipulated at two distinct levels (i.e., high and low for ceiling height, large and small for amount of visible space). Encoding this study-driven parametrization into predictive models will not generalize to new spatial contexts. Instead, we treat these spatial variations as stimulus sampling to enhance external validity (Wells and Windschitl, 1999; Jun et al., 2022).
Formally, we extracted features related to listeners, previous speakers, the dyadic relationship between listeners and speakers, their relationship with the group, and verbal behaviors. We further encoded features related to two users: a user of interest which we refer to as the main user, and the previous speaker, which we refer to as the reference user. When extracting motion-related features, we examined the tracking data of 1-second windows at moments prior to the start of turn-transition behaviors. The 1-second window size is determined based on piloting and aligns with past literature predicting speaking intentions (Chen et al., 2024). Table 2 presents a summary of extracted features.
@@
Dyadic Relationship (Chen et al., 2024; Jokinen et al., 2013; Ishii et al., 2016b; Lee et al., 2023) | Direct visual attention from the main user to the reference user | 3 summary stats. 2 meas. = 6 |
Direct visual attention from the reference user to the main user | 3 summary stats. 2 meas. = 6 | |
@@
Direct visual attention from a user to all remaining users, where the user is either the main user or reference user | 3 summary stats. 2 meas. 2 users = 12 |
Direct visual attention from all remaining users to a user, where the user is either the main user or reference user | 3 summary stats. 2 meas. 2 users = 12 | |
@@ 3.3.1. Speech-Related Features
The first feature group is related to speech behaviors, as they were shown to be predictive of who the next speaker is (Parker, 1988). Following Parker (Parker, 1988), we encoded speech sequences leading up to the start of the examined window. For this, we begin at the timestamp at the start of the examined window, and trace backwards each time the main speaker changes, at which point we increment the turn index by 1. If the main speaker for a given turn is the main user, we assign the feature value at the corresponding turn index “u”. If we encounter a new speaker at a preceding turn, we assign the feature value a new symbol representing that user (e.g., “a”, “b”, “c”). For turns with a speaker who we had already created a symbol for, we assigned their symbol to the feature value. In our setup, we encoded 10 preceding speaker indices, assigning “NA” when there were no more changes in main speakers preceding the last encoded turn. In practice, we processed the speech sequence features at each turn index using one-hot encoding.
We introduced three additional features to describe the main user’s verbal behaviors: (1) the number of speaking turns before the main speaker’s last speech event, (2) the duration of time before the end of the main user’s last speech event, and (3) whether the main user has spoken before during the current session. Deriving these features for the reference user was unnecessary since that user is typically set to the previous speaker.
@@ 3.3.2. Individual- and Group- Level Features
Drawing from works showing correlations between speech behaviors and personalities and group compositions (Ramsay, 1968; Ivanov et al., 2011; Aran and Gatica-Perez, 2013; Mast, 2002; Hadley et al., 2021), we formulated features based on individual and group characteristics. Specifically, we encoded the Big-5 personality traits (Gosling et al., 2003) of the main and reference users, and group average. The authors of the dataset measured Big-5 personality using the ten item personality measure (TIPI) (Gosling et al., 2003), with each personality trait ranging between 1–7 and calculated as the mean of two items.